OpenStudy (anonymous):

Solve for x: e^(2x)-2e^(x)-3=0 So completely lost. Hate natural logarithms. Explanation please?

OpenStudy (anonymous):

@SarahEZZMcK

hero (hero):

Hint: rewrite as (e^x)^2 - 2e^x - 3 = 0 Then substitute y for e^x: y^2 - 2y - 3 = 0 Then factor

OpenStudy (anonymous):

Then solve for y and plug that answer into y=e^x?

hero (hero):

Solve for y, then replace y with e^x. Afterwards, solve for x.

OpenStudy (anonymous):

Factoring leads to (y-3) and (y+1) sooo..... (e^x-3)(e^x+1)=0 ...... now it looks like I need to put this into logarithmic form.... I think?

OpenStudy (anonymous):

no!! hold on....

OpenStudy (anonymous):

e^x=3 and e^x=-1 soooo....

hero (hero):

Solve for x if possible.

OpenStudy (anonymous):

I just revert this into ln form... so... because ln(-1) isn't possible.... the only possible answer is x=ln(3) ?

OpenStudy (anonymous):

WOOHOO!! :D Thank you, sir!